OpenStudy (anonymous):

What is the equation of the line that has a slope of 3 and goes through point (4, 7)?

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] Use this it is the Point-slope formula Substitute the follwing values into the formula m=slope so 3 4=x1 7=y1 (x1,y1)

OpenStudy (anonymous):

y-7=3(x-4) is that all?

pooja195 (pooja195):

Do they want slope -intercept form?

OpenStudy (anonymous):

it just says write an equation, it doesn't specify what type.

pooja195 (pooja195):

hmm @Loser66 @perl

OpenStudy (loser66):

just simplify more

pooja195 (pooja195):

So slope intercept form.... y-7=3(x-4) \[\huge~\rm3(x-4) \] distribute that

OpenStudy (anonymous):

y-7=3x-12 y=3x+19 I have no idea if i did that right

pooja195 (pooja195):

\[\huge~\rm~3 \times-4=??\]

OpenStudy (anonymous):

OH right y=3x+5?

pooja195 (pooja195):

not quite :( answer my question because that will get you the right answer u are very close

pooja195 (pooja195):

-4 x 3 =?

OpenStudy (anonymous):

-12

pooja195 (pooja195):

good y-7=3x-12 Ok now finish it off :)

OpenStudy (anonymous):

y=3x-5?

pooja195 (pooja195):

yes :)

pooja195 (pooja195):

Good job ^_^

OpenStudy (anonymous):

thank you for putting with my terrible math skills I appreciate it!

pooja195 (pooja195):

haha not bad :) theres always room for improvement if you ever need help feel free to pm or tag me :)